# A Survey on Future Frame Synthesis: Bridging Deterministic and Generative Approaches

### v4

@@ Abstract
Future Frame Synthesis (FFS) aims to enable models to generate sequences of future frames based on existing content. This task has garnered widespread application across various domains. In this paper, we comprehensively survey both historical and contemporary works in this field, encompassing widely used datasets and algorithms. Our survey scrutinizes the challenges and the evolving landscape of FFS within the realm of computer vision. We propose a novel taxonomy centered on the stochastic nature of related algorithms. This taxonomy emphasizes the gradual transition from deterministic to generative synthesis methodologies, highlighting significant advancements and shifts in approach.
@@ 1 Introduction
The goal of the future frame synthesis (FFS) task is to generate future frames based on a sequence of historical frames [90] or just a single context frame [122], with or without additional control signals. It is closely related to low-level computer vision processing techniques, particularly when synthesizing near frames [70, 50]. However, FFS diverges from other low-level tasks by implicitly demanding a more complex understanding of scene dynamics and temporal coherence, which is often characteristic of high-level vision tasks. The challenge lies in designing models that can achieve this balance efficiently, using a moderate number of parameters to minimize inference latency and resource consumption, thereby making FFS suitable for real-world applications. This unique position of FFS demonstrates its integral role in bridging the gap between low-level perception & prediction and high-level understanding & generation within computer vision.
@@
Current future frame synthesis algorithms can generally be divided into two categories. One category entails referencing pixels from the current frames (typically the last observed frame) to construct future frames. However, this group of methods inherently confronts difficulties in modeling objects’ appearance and disappearance (birth and death) within the scene. A shared feature of these methods is that they can make highly accurate predictions in the short term, but their accuracy diminishes over longer periods. This kind of research is also commonly referred to as video prediction [76]. The other category encompasses methods that involve generating new frames from scratch. Although these approaches offer the promise of capturing the birth-and-death phenomena in object dynamics, they predominantly focus on modeling pixel-level distributions. As a result, they often lack an integrated understanding of the underlying real-world context, which is crucial for creative synthesis capabilities.
@@
Before our work, two reviews in 2020 on video prediction [76, 79] cover a lot of early technical content. Recently, there have also been literature reviews on “text-to-video generative model” [69] and “long video generation” [65, 93]. Our survey mainly highlights the latest advancements and the connections between predictive and generative methodologies. We believe that the future of long-term FFS hinges on a synergistic integration of prediction and generation techniques. This unified approach would combine contextual constraints with enhanced comprehension, addressing challenges within a cohesive framework.
@@
We propose a novel taxonomy, visually summarized in Figure 1, focusing on the stochasticity of methods. In Section 4, we introduce deterministic algorithms that aim to perform pixel-level fitting based on deterministic target frames. However, due to pixel-level metrics encouraging models to average over multiple equally probable outcomes, they often produce blurred outputs. In Section 5, we discuss algorithms that attempt to imbue models with the ability to make stochastic predictions in motion. This includes methods that introduce stochastic variables or distributions into deterministic models and approaches that directly utilize probabilistic models. Such algorithms allow models to sample from motion distributions, encouraging the generation of predictions that deviate significantly from the target frame but remain reasonable. Given the limited creativity of existing FFS algorithms for high-resolution natural video data, which struggle with creating future frame sequences containing many birth-and-death phenomena, we introduce the generative FFS task in Section 6, which prioritizes generating reasonable video frame sequences in extended temporal durations over pixel-level accuracy. In addition, some related content is placed in the Appendix, including the application scenarios of FFS and the detailed information the training datasets.
@@ 2.1 Problem definition
Many FFS algorithms also make use of additional information, which can be auxiliary data $A_{t_{1}:t_{2}}$ from videos, such as depth maps, landmarks, bounding boxes and segmentation maps, to help the model understand the video. It can also include human input control signals $C_{t_{2}+1:t_{3}}$, such as text instructions and strokes, with the goal of enabling the model to generate future video frame sequences based on specific future trajectory. Based on this consideration, we can extend the formula of the FFS task to a more comprehensive version as Eq. 2:
@@
2.2 Challenges
@@
There exist longstanding challenges in FFS field, including the need for algorithms that balance low-level pixel processing with high-level scene dynamics understanding, the inadequacy of perceptual and stochastic evaluation metrics, the difficulty in achieving long-term synthesis, and limited high-resolution dataset quality for stochastic motion and birth-and-death phenomena. This section provides an overview of challenges, emphasizing the ongoing need for research and development to enhance the accuracy, reliability, and adaptability of FFS algorithms.
High-level metrics.
@@
The low-level metrics such as Peak Signal-to-Noise Ratio (PSNR) and Structural Similarity Index (SSIM), can only assess the pixel-wise accuracy of predicted pixels. To chasing these metrics, researchers typically train models utilizing either the $l_{1}$ or $l_{2}$ loss function of pixel space. The models are prone to average over multiple plausible outcomes, often resulting in blurry predictions know as the perception-distortion tradeoff [10]. Researchers are increasingly looking towards alternatives, such as perceptual metrics (DeePSiM [29], LPIPS [130]), and stochastic metrics (IS [82], FID [46]). These metrics may better align with human perception. At present, reseachers generally use Imagenet classifiers as feature extractors [55, 61] in comparing the generated results with the ground truth to optimize for perception metrics.
@@
While short-term video prediction has seen significant advancements, accurately synthesizing events over extended time horizons remains challenging because of long-term dependencies and complex interactions between objects in dynamic scenes. Due to the inadequacy of model scale in forming a comprehensive understanding of the real world, most existing video synthesis models primarily model changes in pixel distribution. When faced with natural videos spanning extended periods, these models struggle to accurately predict object movements while maintaining visual quality. A promising approach for enhancing long-term synthesis capabilities is the incorporation of high-level structure information [99]. Understanding and utilizing higher-order information can enable the model to retain more key details, thereby achieving internal consistency within the video over longer time scales.
@@
The interplay between data volume and the complexity of models jointly determines the upper bounds of an algorithm’s performance. Despite the vastness of video data available on the internet, the scarcity of high-quality video datasets suitable for video synthesis remains a limiting factor, and existing datasets also pose various challenges, such as simplistic data distribution, low resolution and small motion scales. These issues make it difficult for video synthesis models to handle high-resolution content and large motion scales, thus restricting their practical utility to diverse and unseen scenarios. Achieving high-resolution video synthesis is a complex task demanding significant computational resources [9]. The intricacies involved make real-time applications challenging due to the substantial computational burden.
3 Datasets
The advancement in video synthesis models is greatly dependent on the diversity, quality, and characteristics of training datasets. A general observation is the varying suitability of datasets based on their dimensionality and size, where lower-dimensional datasets, typically with smaller data sizes, may suffer from limited generalizability. In contrast, higher-dimensional datasets provide a broader range of data, contributing to stronger generalization capabilities in the models. Our Appendix offers an overview of the most widely used datasets [83, 26, 60, 89, 54, 36, 132, 88, 57, 90, 22, 1, 31, 77, 38, 39, 24, 23, 121, 6, 87, 126, 123, 12, 17, 15, 109] in video synthesis, highlighting their amounts of data and additional supervisory modalities, thereby providing a comprehensive picture of the current dataset landscape in this field.
@@ 4.1 Raw pixel space
In short-term FFS, methods in raw pixel space have achieved good results.
Recurrent networks.
@@
The exploration of recurrent neural networks in video synthesis is pioneered by PredNet [71], which draws inspiration from predictive coding in neuroscience and employs a recurrent convolutional network for the effective processing of video features. Building on this, PredRNN [110] introduces significant enhancements by modifying Long Short-term Memory (LSTM) with a dual memory structure, aiming for enhanced spatiotemporal modeling. Despite its advancements, it encounters challenges with gradient vanishing in video synthesis tasks. Addressing such limitations, ConvLSTM [85] emerges as a pivotal model, ingeniously integrating LSTM with Convolutional Neural Network (CNN) to proficiently capture motion and spatiotemporal dynamics, a development that has significantly influenced subsequent video synthesis models. Further advancing the field, E3d-LSTM [108] innovatively incorporates 3D convolutions into RNNs and introduces a gate-controlled self-attention module, thereby markedly improving long-term synthesis capabilities. MSPred [97] proposes a hierarchical convolutional and recurrent network operating at multiple temporal frequencies to predict future video frames, as well as other representations such as poses or semantics.
Convolutional networks.
@@
CNNs have been instrumental in the evolution of video synthesis technology. Beginning with GDL [74], the field has seen significant advancements. Following this, PredCNN [120] establishes a new benchmark by outperforming its predecessor PredRNN [110] across various datasets. SDC-Net [80] introduces a novel approach by utilizing a high-resolution video frame synthesis technique that effectively leverages past frames and optical flow. Building on these innovations, the recent introduction of SimVP [35] marks another milestone. This approach revisits the advancements made by ViT [28] and introduces a simplified CNN network, demonstrating that such a configuration can achieve comparable performance in video synthesis. This progression underscores the rapid advancements in convolutional network applications in video synthesis, highlighting a continuous trend toward more efficient and effective models.
Optical-flow-based synthesis.
@@
The optical flow describes the motion of pixels between frames and can be used in flowing pixels of the current frame to synthesize near future frames [70]. Focusing on the enhancement of synthesis quality, FVS [116] utilizes a comprehensive approach by incorporating supplementary information such as semantic maps, instance maps, and optical flow from input frame sequences. This method, while effective, introduces challenges due to increased data modalities and computational demands. OPT [117] estimates the optical flow of video motions in an optimization manner. By continuously refining the current optical flow estimation, the image quality of the next frame can be significantly improved. DMVFN [50] enhances the dense voxel flow [70] estimation by dynamically changing network architecture based on motion magnitude. The study of optical flow estimation is a hot topic [94, 51, 91, 27] in itself. We believe that the near future frames can be synthesized increasingly better under the constantly improving optical flow method. And it is still a challenge to integrate optical flow methods into the goal of improving image synthesis quality [121, 52] and long-term video generation [66].

### v5

@@ Abstract
Future Frame Synthesis (FFS) aims to enable models to generate sequences of future frames based on existing content. This survey comprehensively reviews historical and contemporary works in FFS, including widely used datasets and algorithms. It scrutinizes the challenges and the evolving landscape of FFS within computer vision, with a focus on the transition from deterministic to generative synthesis methodologies. Our taxonomy highlights the significant advancements and shifts in approach, underscoring the growing importance of generative models in achieving realistic and diverse future frame predictions.
@@ 1 Introduction
The goal of the future frame synthesis (FFS) task is to generate future frames based on a sequence of historical frames (Srivastava et al., 2015) or just a single context frame (Xue et al., 2016), with or without additional control signals. The learning objective of this FFS is also considered to be key to building a world model (Ha & Schmidhuber, 2018; Hafner et al., 2023b). FFS is closely related to low-level computer vision processing techniques, particularly when synthesizing near frames (Liu et al., 2017; Wu et al., 2022b; Hu et al., 2023b). However, FFS diverges from other low-level tasks by implicitly demanding a more complex understanding of scene dynamics and temporal coherence, which is often characteristic of high-level vision tasks. The challenge lies in designing models that can achieve this balance efficiently, using a moderate number of parameters to minimize inference latency and resource consumption, thereby making FFS suitable for real-world applications. This unique position of FFS demonstrates its integral role in bridging the gap between low-level perception & prediction and high-level understanding & generation within computer vision.
@@
Current FFS algorithms can generally be divided into two categories. One category entails referencing pixels from the current frames (typically the last observed frame) to construct future frames. However, this group of methods inherently confronts difficulties in modeling objects’ appearance and disappearance (birth and death) within the scene. A shared feature of these methods is that they can make highly accurate predictions in the short term, but their accuracy diminishes over longer periods. This kind of research is also commonly referred to as video prediction (Oprea et al., 2020). The other category encompasses methods that involve generating new frames from scratch. Although these approaches offer the promise of capturing the birth-and-death phenomena in object dynamics, they predominantly focus on modeling pixel-level distributions. As a result, they often lack an integrated understanding of the underlying real-world context, which is crucial for creative synthesis capabilities.
@@
Before our work, two reviews in 2020 on video prediction (Oprea et al., 2020; Rasouli, 2020b) cover a lot of early technical content. Recently, there have also been literature reviews on "text-to-video generative model" (Liu et al., 2024) and "long video generation" (Li et al., 2024; Sun et al., 2024b). Our survey mainly highlights the latest advancements and the connections between predictive and generative methodologies. We believe that the future of long-term FFS hinges on a synergistic integration of prediction and generation techniques. This unified approach would combine contextual constraints with enhanced comprehension, addressing challenges within a cohesive framework.
@@
As background, we will introduce the problem, the dataset, and the fundamental challenges faced in Section 2. Our taxonomy focuses on the stochasticity of methods. In Section 3, we introduce deterministic algorithms that aim to perform pixel-level fitting based on deterministic target frames. However, due to pixel-level metrics encouraging models to average over multiple equally probable outcomes, they often produce blurred outputs. In Section 4, we discuss algorithms that attempt to imbue models with the ability to make stochastic predictions in motion. This includes methods that introduce stochastic variables or distributions into deterministic models and approaches that directly utilize probabilistic models. Such algorithms allow models to sample from motion distributions, encouraging the generation of predictions that deviate significantly from the target frame but remain reasonable. Given the limited creativity of existing FFS algorithms for high-resolution natural video data, which struggle with creating future frame sequences containing many birth-and-death phenomena, we introduce the generative FFS task in Section 5, which prioritizes generating reasonable video frame sequences in extended temporal durations over pixel-level accuracy. In Section 6, we explore the diverse applications of FFS across various domains, highlighting its significance in autonomous driving, robot navigation, the cinema industry, meteorology, and anomaly detection. These applications underscore the importance and potential of FFS as a tool for understanding and interacting with the world around us. In Section 7, we provide an overview of previous surveys on video prediction and video diffusion models. We also clarify the focus of our survey, which comprehensively reviews historical and contemporary works in FFS, focusing on the transition from deterministic to generative synthesis methodologies. It emphasizes the growing importance of generative models in achieving realistic and diverse future frame predictions, highlighting significant advancements and shifts in approach.
@@ 2.1 Problem Definition
Many FFS algorithms also make use of additional information, which can be auxiliary data $A_{t_{1}:t_{2}}$ from videos, such as depth maps, landmarks, bounding boxes, and segmentation maps, to help the model understand the video. It can also include human input control signals $C_{t_{2}+1:t_{3}}$, such as text instructions and strokes, with the goal of enabling the model to generate future video frame sequences based on specific future trajectories. Based on this consideration, we can extend the formula of the FFS task to a more comprehensive version as Eq. 2:
@@
Dataset Category # Videos # Clip Frames Resolution Extra Annotations
KTH Action (Schuldt et al., 2004) Human 2, 391 95 160 $\times$ 120 Class
Caltech Pedestrian (Dollar et al., 2011) Human 137 1, 824 640 $\times$ 480 Bounding Box
HMDB51 (Kuehne et al., 2011) Human 6, 766 93 414 $\times$ 404$\dagger$ Class
UCF101 (Soomro et al., 2012) Human 13, 320 187 320 $\times$ 240 Class
J-HMDB (Jhuang et al., 2013) Human 928 34 320 $\times$ 240 OF, Ins, HJ, Class
KITTI (Geiger et al., 2013) Traffic 151 323 1242 $\times$ 375 OF, BBox, Sem, Ins, Depth
Penn Action (Zhang et al., 2013) Human 2, 326 70 480 $\times$ 270$\dagger$ Human Joint, Class
SJTU 4K (Song et al., 2013) General 15 300 3840 $\times$ 2160 -
Sports-1M (Karpathy et al., 2014) Human 1, 133, 158 variable variable Class
Moving MNIST (Srivastava et al., 2015) Simulation 10, 000 20 64 $\times$ 64 -
Cityscapes (Cordts et al., 2016) Traffic 46 869 2048 $\times$ 1024 Semantic, Instance, Depth
YouTube-8M (Abu-El-Haija et al., 2016) General 8, 200, 000 variable variable Class
Robotic Pushing (Finn et al., 2016) Robot 59, 000 25 640 $\times$ 512 Class
DAVIS17 (Pont-Tuset et al., 2017) General 150 73 3840 $\times$ 2026$\dagger$ Semantic
Something-Something (Goyal et al., 2017) Object 220, 847 45 427 $\times$ 240$\dagger$ Text
ShapeStacks (Groth et al., 2018) Simulation 36, 000 16 224 $\times$ 224 Semantic
SM-MNIST (Denton & Fergus, 2018) Simulation customize customize 64 $\times$ 64 -
D2-City (Che et al., 2019) Traffic 11, 211 750 1080p / 720p BBox
Kinetics-700 (Carreira et al., 2019) Human 650, 000 250 variable Class
RoboNet (Dasari et al., 2019) Robot 161, 000 93 64 $\times$ 48 -
Vimeo-90K (Xue et al., 2019) General 91, 701 7 448 $\times$ 256 -
BDD100K (Yu et al., 2020) Traffic 100, 000 1175 1280 $\times$ 720 BBox, Semantic, Depth
nuScenes (Caesar et al., 2020) Traffic 1, 000 40 1600 $\times$ 900 BBox, Semantic
WebVid (Bain et al., 2021) General 10, 732, 607 449 596 $\times$ 336 Text
X4K1000FPS (Sim et al., 2021) General 4, 408 65 4096 $\times$ 2160 -
SportsSlomo (Chen & Jiang, 2024) Human 130,000 7 $1280\times 720$
InternVideo2 (Wang et al., 2024b) General 2, 000, 000 variable variable Text, Action
OpenDV-YouTube (Yang et al., 2024) Traffic 2, 139 100,000+ variable Text
denotes the mean value. $\dagger$ denotes the median value.
2.2 Datasets
The advancement in video synthesis models is greatly dependent on the diversity, quality, and characteristics of training datasets. A general observation is the varying suitability of datasets based on their dimensionality and size, where lower-dimensional datasets, typically with smaller data sizes, may suffer from limited generalizability. In contrast, higher-dimensional datasets provide a broader range of data, contributing to stronger generalization capabilities in the models. We offer an overview of the most widely used datasets in video synthesis in Table 1, highlighting their amounts of data and additional supervisory modalities, thereby providing a comprehensive picture of the current dataset landscape in this field. In cases where specific dataset details are not reported either in the original paper or the project page, we calculate the mean or median statistics to maintain consistency in our analysis.
Challenges.
1. Unify the organization of image and video data. A substantial amount of computer vision work is conducted on the image modality. Therefore, image datasets may undergo more meticulous curation and possess more annotated data. The most commonly used billions-level image datasets include YFCC100M (Thomee et al., 2016), WIT400M (Radford et al., 2021) and LAION400M (Schuhmann et al., 2021). Given the vast amount of data involved, researchers may need to effectively leverage knowledge from foundational image models. When incorporating video data into training, we may need to filter out low-quality segments and select an appropriate sampling frame rate.
2. Determine the proportion of data from different domains. Computer graphics composite data, 2D anime data, real videos, and videos with special effects can have vastly different appearances. Moreover, we may not be able to unify data from different sources into a specific resolution because videos may have different aspect ratios, and details that only make sense at high resolutions (such as subtitles, and textures). Many frame synthesis methods are sensitive to resolution because of the correlation between resolution and object motion intensity (Sim et al., 2021; Hu et al., 2023b; Yoon et al., 2024).
2.3 Overall Challenges
@@
There exist longstanding challenges in FFS field, including the need for algorithms that balance low-level pixel processing with high-level scene dynamics understanding, the inadequacy of perceptual and stochastic evaluation metrics, the difficulty in achieving long-term synthesis, and limited high-resolution dataset quality for stochastic motion and birth-and-death phenomena. This section provides an overview of challenges.
Evaluation metrics.
@@
The low-level metrics such as Peak Signal-to-Noise Ratio (PSNR) and Structural Similarity Index (SSIM), can only assess the pixel-wise accuracy of predicted pixels. To chase these metrics, researchers typically train models utilizing either the $l_{1}$ or $l_{2}$ loss function of pixel space. The models are prone to average over multiple plausible outcomes, often resulting in blurry predictions, known as the perception-distortion tradeoff (Blau & Michaeli, 2018). They tend to favor blurry predictions that nearly accommodate the ground truth over sharper and more plausible but imperfect generations that do not match the ground truth. Many researchers are increasingly looking towards alternatives, such as perceptual metrics (DeePSiM (Dosovitskiy & Brox, 2016), LPIPS (Zhang et al., 2018)), and stochastic metrics (IS (Salimans et al., 2016), FID (Heusel et al., 2017)). These metrics may better align with human perception. But even classifiers trained on human perception annotations have a relatively low level of agreement with humans in judging image quality (Kumar et al., 2022).
Visual researchers typically evaluate models based on the visual quality of the generated results. However, for many specific applications, we are not sure that visual quality is crucial. For instance, Dreamer-V3 (Hafner et al., 2023a) and VPT (Baker et al., 2022) have successfully built visual models on low-resolution frame sequences. Moreover, most research work on visual representations is built on relatively small image resolutions (Radford et al., 2021; He et al., 2022). We are concerned that an excessive pursuit of visual quality may lead us to favor models that focus on low-level features. In addition to matching human subjective perception, we also need to design metrics that evaluate a model’s ability to capture scene dynamics and temporal variations.
Even with better evaluation metrics, optimizing for them is a challenge. During model training, researchers often use Imagenet classifiers as feature extractors (Johnson et al., 2016; Kumar et al., 2022) in comparing the generated results with the ground truth to optimize for both low-level and high-level features. There are also some GAN-based loss functions that can improve the image quality of the generated results (Huang et al., 2017; Zhang et al., 2020).
@@
While short-term video prediction has seen significant advancements, accurately synthesizing events over extended time horizons remains challenging because of long-term dependencies and complex interactions between objects in dynamic scenes. Iteratively using a short-term video prediction model will quickly lead to degraded results (Wu et al., 2022b; Hu et al., 2023b). Due to the inadequacy of model scale in forming a comprehensive understanding of the real world, most existing video synthesis models primarily model changes in pixel distribution. When faced with natural videos spanning extended periods, these models struggle to accurately predict object movements while maintaining visual quality. A promising approach for enhancing long-term synthesis capabilities is the incorporation of high-level structure information (Villegas et al., 2017). Understanding and utilizing higher-order information can enable the model to retain more key details, thereby achieving internal consistency within the video over longer time scales.
@@
The interplay between data volume and the complexity of models jointly determines the upper bounds of an algorithm’s performance. Despite the vastness of video data available on the internet, the scarcity of high-quality video datasets suitable for video synthesis remains a limiting factor, and existing datasets also pose various challenges, such as simplistic data distribution, low resolution, and small motion scales. These issues make it difficult for video synthesis models to handle high-resolution content and large motion scales, thus restricting their practical utility to diverse and unseen scenarios. Achieving high-resolution video synthesis is a complex task demanding significant computational resources (Blattmann et al., 2023b). The intricacies involved make real-time applications challenging due to the substantial computational burden.
@@ 3.1 Raw Pixel Space
In short-term FFS, methods in raw pixel space have achieved good results. We hope to review existing methods and discuss the challenges faced.
3.1.1 Recurrent networks.
@@
The exploration of recurrent neural networks in video synthesis is pioneered by PredNet (Lotter et al., 2016), which draws inspiration from predictive coding in neuroscience and employs a recurrent convolutional network for the effective processing of video features. Building on this, PredRNN (Wang et al., 2017) introduces significant enhancements by modifying Long Short-term Memory (LSTM) with a dual memory structure, aiming for enhanced spatiotemporal modeling. Despite its advancements, it encounters challenges with gradient vanishing in video synthesis tasks. Addressing such limitations, ConvLSTM (Shi et al., 2015) emerges as a pivotal model, ingeniously integrating LSTM with Convolutional Neural Network (CNN) to proficiently capture motion and spatiotemporal dynamics, a development that has significantly influenced subsequent video synthesis models. Further advancing the field, E3d-LSTM (Wang et al., 2018b) innovatively incorporates 3D convolutions into RNNs and introduces a gate-controlled self-attention module, thereby markedly improving long-term synthesis capabilities. Nonetheless, the increased computational complexity due to 3D convolutions might offset the performance gains in certain applications. MSPred (Villar-Corrales et al., 2022) proposes a hierarchical convolutional and recurrent network operating at multiple temporal frequencies to predict future video frames, as well as other representations such as poses or semantics.
Challenges.
Recurrent networks, despite their effectiveness in capturing temporal dependencies, face several challenges in video prediction tasks. Their sequential nature, which allows them to model frame-by-frame changes, can lead to high computational complexity, especially in high-resolution scenarios. This is evident from the significantly higher FLOPs and lower FPS observed in recurrent-based models compared to their recurrent-free counterparts (Tan et al., 2023). Additionally, recurrent networks are prone to gradient vanishing and exploding issues, which can hinder their ability to learn long-term dependencies (Gao et al., 2022). These challenges underscore the need for alternative approaches that can balance efficiency and performance, such as recurrent-free models, which have shown promising results in various video prediction tasks.
3.1.2 Convolutional networks.
@@
CNNs have been instrumental in the evolution of video synthesis technology. Beginning with GDL (Mathieu et al., 2015), the field has seen significant advancements. Following this, PredCNN (Xu et al., 2018) establishes a new benchmark by outperforming its predecessor PredRNN (Wang et al., 2017) across various datasets. SDC-Net (Reda et al., 2018) introduces a novel approach by utilizing a high-resolution video frame synthesis technique that effectively leverages past frames and optical flow. Building on these innovations, the introduction of SimVP (Gao et al., 2022) marks another milestone. This approach revisits the advancements made by ViT (Dosovitskiy et al., 2021) and introduces a simplified CNN network, demonstrating that such a configuration can achieve comparable performance in video synthesis.
Challenges.
The CNN-based frame synthesis method, although simple to implement and fast, is not suitable for spatially shifting the pixels of the input frames. CAIN (Choi et al., 2020) and FLAVR (Kalluri et al., 2023) introduced channel attention and 3D-UNet in the synthesis of intermediate frames, respectively, but they do not completely replace explicit pixel motion methods such as kernel-based methods and flow-based methods. Moreover, in pursuit of efficiency, most CNN networks used for FFS have a relatively small number of parameters, generally not exceeding 60M (Tan et al., 2023). As a comparison, in order to leverage large datasets, the video diffusion model has been scaled up to over 1.5B parameters (Blattmann et al., 2023a). Scaling up CNN-based models effectively presents a significant challenge. We speculate that short-term prediction models for high-resolution real-time applications and models that attempt to draw knowledge and generation capabilities from large datasets will diverge.
3.1.3 Optical-flow-based synthesis.
@@
The optical flow describes the motion of pixels between frames and can be used in flowing pixels of the current frame to synthesize near future frames (Liu et al., 2017). We consider the flow-based methods extensions of the kernel-based methods (Niklaus et al., 2017), as the kernel-based methods typically constrain the movement of pixels to a relatively small neighborhood (Cheng & Chen, 2021). Focusing on the enhancement of synthesis quality, FVS (Wu et al., 2020) utilizes a comprehensive approach by incorporating supplementary information such as semantic maps, instance maps, and optical flow from input frame sequences. This method, while effective, introduces challenges due to increased data modalities and computational demands. OPT (Wu et al., 2022b) estimates the optical flow of video motions in an optimization manner. By continuously refining the current optical flow estimation, the image quality of the next frame can be significantly improved. This approach effectively leverages the knowledge from off-the-shelf optical flow models (Teed & Deng, 2020) and frame interpolation models (Jiang et al., 2018; Huang et al., 2022b). Although training is omitted, the iterative optimization process during each inference requires substantial computation. DMVFN (Hu et al., 2023b) enhances the dense voxel flow (Liu et al., 2017) estimation by dynamically changing network architecture based on motion magnitude. DMVFN confirms the importance of a coarse-to-fine, multi-scale approach in solving short-term flows.
Challenges.
The study of optical flow estimation is a hot topic in itself (Teed & Deng, 2020; Huang et al., 2022a; Sun et al., 2022; Dong & Fu, 2024). However, mainstream optical flow models trained on synthetic data with strong augmentations may differ significantly from the scenarios focused on by FFS. Moreover, the learning objective of these models is not to provide a flow that is suitable for moving pixels and synthesizing high-quality images.  Xue et al. (2019) point out that for different downstream tasks, we need to fine-tune or even train a flow estimation network from scratch. Higher-performance optical flow networks may even lead to worse image synthesis effects because they may focus on ambiguous areas such as occlusions and have insufficient resolution (Niklaus & Liu, 2020; Huang et al., 2022b). In real-world scenarios, it is also challenging to obtain optical flow labels directly.
We believe that the near future frames can be synthesized increasingly better under the constantly improving optical flow method. And it is still a challenge to integrate optical flow methods into the goal of long-term video generation (Liang et al., 2023). Optical flow may only be used to predict pixel movement over a very short period of time and cannot help generate new video content.
3.1.4 Transformers.
After ViT’s groundbreaking design of a pure transformer applied directly to sequences of image patches Dosovitskiy et al. (2021), the application of transformers to frame synthesis has become a hot topic. Video frame interpolation is a task very closely related to FFS (Liu et al., 2017). Both  Shi et al. (2022) and  Lu et al. (2022) propose transformer-based video interpolation frameworks to overcome the limitations of traditional CNNs, leveraging self-attention mechanisms to capture long-range dependencies and enhance content-awareness. They introduce innovative strategies, such as local attention in the spatial-temporal domain and cross-scale window-based attention, to improve performance and handle large motions effectively.  Ye & Bilodeau (2023) present an efficient transformer model for video prediction, leveraging a novel local spatial-temporal separation attention mechanism, and compares three variants (fully auto-regressive, partially auto-regressive, and non-autoregressive) to achieve optimal performance and reduced complexity. There are still many ongoing studies aimed at advancing the use of transformers to address inter-frame dynamics with varying motion amplitudes and high-resolution synthesis issues (Park et al., 2023; Zhang et al., 2024).
Challenges.
Many researchers believe that transformers have an advantage on large datasets (Zhai et al., 2022; Smith et al., 2023). However, most of these studies focus on high-level vision tasks. Moreover, the success of transformers on many tasks relies on fully leveraging the capabilities of foundational models. In image synthesis tasks, we are not yet certain about the best practices (Li et al., 2021). Leveraging the experience of large language models (LLMs) and similar architectures could be a promising direction, which we will discuss in subsequent Section 5.2.